# About carrier’s self-trapping and dynamical Rashba splitting in the two-dimensional hybrid perovskite (BA)2(MA)2Pb3I10

### v2

@@ Abstract
Time- and Angle-Resolved Photoelectron Spectroscopy (tr-ARPES) is employed to monitor photoexcited electrons in the two-dimensional hybrid perovskite (BA)2(MA)2Pb3I10. Photoelectron intensity maps are in good agreement with ab-initio calculations of the band structure. The effective mass is and for holes and electrons, respectively. In the photoexcited state, spin-orbit splitting of the conduction band cannot be resolved. This sets the upper bound of photoinduced Rashba coupling to eVÅ. The correlated electron-hole plasma evolves in Wannier excitons with Bohr radius of 2.5 nm, while no sign of self-trapping in small polarons is found within the investigated time window of up to 120 ps following photoexcitation.
@@
We performed PBE-DFT calculations for the experimentally detetermined structure [10] of (BA)2(MA)2Pb3I10. The computations employed a plane-wave energy cutoff of 40 Hartree and a 4×2×4 Monkhorst-Pack k-point grid. Spin–orbit coupling was included via fully relativistic pseudopotentials [13, 14]. Figure 1c) shows the electronic band structure calculated along the high-symmetry path X(1/2,0,0) – (0,0,0) – R(1/2,0,1/2). The PBE-DFT band gap at the -point is found to be 0.52 eV, significantly lower than the experimental value of 2.1 eV [10]. This discrepancy is a well-known limitation of DFT calculations, which consistently underestimate band gap values. Theoretical estimate of effective masses are obtained by fitting the band dispersion along the path close to the band edges, resulting in and for holes and electrons, respectively. The electronic bands exhibit complex Spin-Orbit splitting (SO), that varies significantly with the crystallographic direction. As shown in Fig. 1c), the SO is negligible along the direction while attains the maximal value meV along the direction. Such anisotropic SO [15] mainly originates from in-plane asymmetric tilting of adjacent metal halide octahedra. A small wavevector expansion along the direction theoretically provides a Rashba parameter eVÅ for the electronic states of the conduction band.
@@
Bulk single crystals of (BA)2(MA)2Pb3I10 were grown by the temperature gradient method [16]. ARPES maps have been acquired on the TEMPO beamlime of Synchrotron Soleil, with a hemispherical analyzer MBS A-1 equipped with vacuum tunable lens axis. The photon beam is centered at 55 eV, horizontally polarized and focused down to 100 300 microns. Single crystals of (BA)2(MA)2Pb3I10 have been cleaved in ultra-high vacuum conditions and measured at 80 K. The natural cleavage plane lies in the buffer layer of the BA ligands, exposing a highly passivated surface. However two-dimensional hybrid perovskites can be easily damaged by XUV photons, so that no ARPES data have ever been reported yet. In this work, we have successfully overcome this limitation by implementing a rapid rastering acquisition mode for ARPES and tr-ARPES, enabling the direct measurement of the band structure in these materials.
@@
As shown in Fig. 2a) the electronic states disperse along the direction, reaching the VBM at . By integrating the intensity map around the VBM, we obtain a Momentum Distribution Curve (MDC) exhibiting peaks at . Also note that a weak, folded replica is generated at by octahedral distortions. Instead, by integrating the intensity map around we obtain an Energy Distribution Curve (EDC) peaking at roughly -2.1 eV (we set the zero energy of ARPES maps at the conduction band minimum). Of the 400 meV peak broadening, a contribution of meV can be ascribed, via ab-initio simulations, to the intrinsic electron-phonon linewidth [17], while the remaining 300 meV contribution is likely of inhomogeneous and extrinsic origin.
@@
The dynamics of electronic states in the conduction band is obtained via a two photon photoemission process: the probe beam with photon energy at eV is horizontally polarized and focused down to 100 100 microns; the pump beam with photon energy of 3.1 eV has incident fluence of 14 J/cm2. The cross correlation between pump and probe has a Full Width Half Maximum of roughly 120 fs [21]. Single crystals of (BA)2(MA)2Pb3I10 are cleaved at 300 K in ultra-high vacuum conditions and measured at temperature below 140 K. Also in this case, the sample has been rastered during data acquisition.
@@
After 0.4 ps (see Fig. 3d) the electrons have reached the CBM, leading to a correlated electron plasma dressed by strong Coulomb interactions with VB holes. As can be seen in Fig. 3d-f, the ARPES intensity decreases substantially (more precisely a factor 4) when moving from a delay of 0.4 ps to 120 ps. Unlike in the case of 3D perovskites [22], the reduced signal intensity is entirely due to carrier recombination. The reason is that the BA ligand hinders the interlayer hopping of electrons, preventing the carrier diffusion from the surface to the bulk. Notice that the ARPES signal of Fig. 3d-f is sharply peaked at the center of the Brillouin Zone and shifts to lower energy by about 70 meV at delay times of 120 ps. Such an energy shift is comparable to the binding energy of exciton [32], revealing the evolution from a correlated electron-hole plasma (at 0.4 ps) to a low density exciton gas (at 120 ps). Recombination does indeed reduce the photo excited density of electrons and holes, until the many-particle screening becomes ineffective.
@@
Since our excitation is not resonant, the ARPES intensity maps in Fig. 3d–f do not exhibit the negative exciton dispersion that has been reported in transition metal dichalcogenides [26, 27]. This Floquet dressing is not always observed [28, 29, 30], because is enhanced by resonant excitation conditions and may depend on specific material system. Theoretical studies also indicate that under non-resonant excitation, the signal is dominated the thermal and incoherent population of excitons with finite center-of-mass momentum [31].
@@
Further conclusions can be drawn from the distribution of spectral weight in wavevector space. Figure 4a shows the MDCs extracted by integrating the ARPES intensity map in an energy interval of 0.1 eV around the CBM, at a delay time of 0.4 ps (dark circles) and 120 ps (red triangles). The two MDCs can be nicely superimposed by applying a relative multiplication factor. They both peak at and can be fitted by the function . Here is the 2D Fourier transform of , namely the model function of a two dimensional exciton with Bohr radius . Our best estimate nm gives nm, in agreement with the spectroscopy of Landau levels at high magnetic fields [32].
@@
The area covered by (i.e. nm2) contains 35 unit cells, therefore excluding the self-trapping of excitons in polaronic cages having the size of inter-ionic distance. This result is important for the ongoing debate about whether excitons are trapped extrinsically or intrinsically in these kinds of materials [33]. We conclude that the broadband emission spectrum reported in the subgap emission observed in (BA)2(MA)n-1PbnI3n+1 [34] are likely due to exciton trapping by extrinsic defects.
@@
The last topic to discuss is dynamical Spin Orbit splitting (SO). Transient polarization of electronic states with eVÅ have been proposed by theoretical simulations [6] and claimed by the spectral analysis of two photon absorption [35]. The proposed conjecture is that excited electrons interacts with the ionic lattice, forming fluctuating ferroelectric domains which could locally increase the SO. As a consequence, we checked if photo-excitation could indeed substantially enhance the Rashba coupling parameter. Figure 4b plots the fitting curve of the experimental MDCs (black line), and MDC model assuming eVÅ (pink shadow area). The absence of a double peak in the experimental MDC let us conclude that eVÅ. Same result is obtained analyzing data acquired along the direction. This upper bound of is consistent with our ab-initio calculations and it excludes that fluctuating ferroelectric domains can generate a dynamical Rashba splitting 10 times larger than the static value.
@@
In conclusion, ARPES and TR-ARPES experiments enable us to visualise the dispersion of electronic states within the valence and conduction bands. This challenging measurement has only been possible here thanks to sample rastering during data acquisition. Our intensity maps show experimental effective masses in good agreement with first-principles calculations. First, excited electrons relax to the bottom of the conduction band, and then they form bound states with the underlying holes. The wavevector tomography of the resulting wavefunction is compatible with Wannier excitons with a Bohr radius of 2.5 nm. No evidence of self-trapping is observed until 120 ps after photoexcitation. Moreover, a thorough analysis of the momentum distribution curve also excludes that fluctuating ferroelectric domains can give rise to dynamic Rashba splitting of very large amplitude.
@@
We acknowledge financial support of the 2D-HYPE project from the Agence Nationale de la Recherche (ANR, Nr. ANR 21-CE30-0059), of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, Nr. 490867834), of the MINTOAURE project (ANR A-22-PETA-0015) and of the Région Ile de France via (FemtoARPES2.0 project N. EX079194). Weiyan Qi thanks the China Scholarship Council (CSC).

### v3

@@ Abstract
Time- and Angle-Resolved Photoelectron Spectroscopy (tr-ARPES) is employed to monitor photoexcited electrons in the two-dimensional hybrid perovskite (BA)2(MA)2Pb3I10. Photoelectron intensity maps are in good agreement with ab-initio calculations of the band structure. The effective mass is and for holes and electrons, respectively. In the photoexcited state, spin-orbit splitting of the conduction band cannot be resolved. This sets the upper bound of photoinduced Rashba coupling to eVÅ. The correlated electron-hole plasma evolves in Wannier excitons with Bohr radius of 2.8 nm, while no sign of self-trapping in small polarons is found within the investigated time window of up to 120 ps following photoexcitation.
@@
We performed PBE-DFT calculations for the experimentally detetermined structure [10] of (BA)2(MA)2Pb3I10. The computations employed a plane-wave energy cutoff of 40 Hartree and a 4×2×4 Monkhorst-Pack k-point grid. Spin–orbit coupling was included via fully relativistic pseudopotentials [13, 14]. Figure 1c) shows the electronic band structure calculated along the high-symmetry path X(1/2,0,0) – (0,0,0) – S(1/2,0,1/2) and in the out of plan direction . The PBE-DFT band gap at the -point is found to be 0.52 eV, significantly lower than the experimental value of 2.1 eV [10]. This discrepancy is a well-known limitation of DFT calculations, which consistently underestimate band gap values. Theoretical estimate of effective masses are obtained by fitting the band dispersion along the path close to the band edges, resulting in and for holes and electrons, respectively. These values are somewhat larger than reported PBEsol-DFT [10] calculations of the in-plane averaged effective masses. We have performed a PBEsol-DFT calculations and have found a % reduction of the hole and electron effective masses giving a good agreement with Stoumpos et al. [10] (see Fig. 5 in the supplementay information file). The electronic bands exhibit complex Spin-Orbit splitting (SO), that varies significantly with the crystallographic direction. As shown in Fig. 1c), the SO is negligible along the direction while attains the maximal value meV along the direction. Such anisotropic SO [15] mainly originates from in-plane asymmetric tilting of adjacent metal halide octahedra. A small wavevector expansion along the direction theoretically provides a Rashba parameter eVÅ for the electronic states of the conduction band. This value is consistent with the equilibrium SO splitting calculated by Jana et al. for similar compounds [15] and is one order of magnitude smaller than the SO splitting predicted by Etienne et al. [6] for a fluctuating lattice.
@@
Bulk single crystals of (BA)2(MA)2Pb3I10 were grown by the temperature gradient method [16]. ARPES maps have been acquired on the TEMPO beamlime of Synchrotron Soleil, with a hemispherical analyzer MBS A-1 equipped with vacuum tunable lens axis. The photon beam is centered at 55 eV, horizontally polarized and focused down to 100 300 microns. Single crystals of (BA)2(MA)2Pb3I10 have been cleaved in ultra-high vacuum conditions and measured at 80 K. The natural cleavage plane lies in the buffer layer of the BA ligands, exposing a highly passivated surface. However two-dimensional hybrid perovskites can be easily damaged by XUV photons, so that no ARPES data have ever been reported yet. In this work, we have successfully overcome this limitation by implementing a rapid rastering acquisition mode for ARPES and tr-ARPES, enabling the direct measurement of the band structure in these materials (details are given in the supplementary information file).
@@
As shown in Fig. 2a) the electronic states disperse along the direction, reaching the VBM at . By integrating the intensity map around the VBM, we obtain a Momentum Distribution Curve (MDC) exhibiting peaks at . Also note that a weak, folded replica is generated at by octahedral distortions. Instead, by integrating the intensity map around we obtain an Energy Distribution Curve (EDC). Of the 400 meV peak broadening, a contribution of meV can be ascribed, via ab-initio simulations, to the intrinsic electron-phonon linewidth [17], while the remaining 300-meV contribution is likely to be dominated by inhomogeneous and extrinsic factors.
@@
The dynamics of electronic states in the conduction band is obtained via a two photon process. The probe beam, is horizontally polarized, has photon energy of eV and is focused down to 100 100 microns. The pump beam has photon energy of 3.1 eV and has incident fluence of 10 J/cm2, corresponding to photoexcitation density cm-2. The cross correlation between pump and probe has a Full Width Half Maximum of roughly 120 fs [21]. Single crystals of (BA)2(MA)2Pb3I10 are cleaved at 300 K in ultra-high vacuum conditions and measured at temperature below 140 K. Also in this case, the sample has been rastered during data acquisition (see also supplementary information file).
@@
After 0.4 ps (see Fig. 3d) the electrons have relaxed, either forming a ionized carriers near to the conduction band minimum [24], or already forming hot excitons. The second hypothesis is more likely as long as the photoexcitation density is not large enough to destabilize the bound el-h states. To prove that excitons are stable at our photoexcitation density, we monitored the tr-ARPES intensity while scanning a pump beam across the optical bandgap. Figure 4 in the supplementary information file shows an exciton resonance peaking at an optical gap of 2.1 eV. The resonance signal corresponds to the photoexcitation density , confirming the possible formation of excitons also for the above gap pumping and at early delay time (0.4 ps). As can be seen in Fig. 3d-f, the ARPES intensity decreases substantially (more precisely a factor 3) when moving from a delay of 0.4 ps to 120 ps. Unlike in the case of 3D perovskites [22], the reduced signal intensity is entirely due to carrier recombination. The reason is that the BA ligand hinders the interlayer hopping of electrons, preventing the carrier diffusion from the surface to the bulk. Notice that the ARPES signal of Fig. 3d-f is sharply peaked at the center of the Brillouin Zone and shifts to lower energy by about 60 meV at delay times of 120 ps (see also Fig. 3 in supplementary information file). Such an energy shift suggests the evolution from dense and hot exciton ensemble (at 0.4 ps) to a low density exciton gas (at 120 ps).
@@
Since our excitation is not resonant, the ARPES intensity maps in Fig. 3d–f do not exhibit the negative exciton dispersion that has been reported in transition metal dichalcogenides [27, 28]. This Floquet dressing is not always observed [29, 30, 31], because is enhanced by resonant excitation conditions and may depend on specific material system. Theoretical studies also indicate that, in the case of excitation above the gap, the signal is dominated the thermal and incoherent population of excitons with finite center-of-mass momentum [32].
@@
Further conclusions can be drawn from the distribution of spectral weight in wavevector space. Figure 4a shows the MDCs extracted by integrating the ARPES intensity map in an energy interval of 0.1 eV around the CBM, at a delay time of 0.4 ps (dark circles) and 120 ps (red triangles). The two MDCs can be nicely superimposed by applying a relative multiplication factor. They both peak at and can be fitted by the function convoluted by the wavevector resolution of the setup. Here is the 2D Fourier transform of , namely the model function of a two dimensional exciton with Bohr radius . The simple 2D hydrogenic model yields nm when considering the exciton effective mass , together with an effective in-plane dielectric constant . The fitting of the experimental data results in nm, which indeed falls within the expected range. Furthermore, the exciton extension nm is in good agreement with the spectroscopy of Landau levels at high magnetic fields [33].
@@
The area covered by (i.e. nm2) contains 45 unit cells, therefore excluding the self-trapping of excitons in polaronic cages having the size of inter-ionic distance. This result is important for the ongoing debate about whether excitons are trapped extrinsically or intrinsically in these kinds of materials [34]. Indeed, the formation of localized, Frenkel-like, excitons would lead to a dispersionless state across the wave-vector window considered here. We conclude that the subgap emission observed in (BA)2(MA)n-1PbnI3n+1 [35] are likely due to exciton trapping by a small concentration of extrinsic defects, which may be visible in integrated photo-luminescence but is irrelevant in a tr-ARPES experiment.
@@
The last topic to discuss is dynamical Spin Orbit splitting (SO). Transient spin polarization of electronic states with eVÅ have been proposed by theoretical simulations [6] and claimed by the spectral analysis of two photon absorption [36]. The proposed conjecture is that ionic lattice motion and dynamical formation of fluctuating ferroelectric domains could locally increase the SO splitting. This effect could be even boosted if excited electrons would self trap in polaronic states. However, designing experiments that could ambiguously estimate the dynamical SO is especially challenging. For example, the difference in threshold energy between one-photon and two-photon absorption could be dominated by the optical selection rules of the exciton lines rather than SO splitting. The tr-ARPES is a far more direct approach for monitoring the potential presence of dynamical Rashba spitting. Figure 4b plots the fitting curve of the experimental MDCs (black line), and the MDC model assuming eVÅ (pink shadow area). The absence of a double peak in the experimental MDC let us conclude that eVÅ. Same result is obtained analyzing data acquired along the direction. This upper bound of is consistent with our ab-initio calculations and it excludes that fluctuating ferroelectric domains can generate a dynamical Rashba splitting 10 times larger than the frozen lattice value.
@@
In conclusion, ARPES and TR-ARPES experiments on 2D hybrid Perovkistes enable us to visualise the dispersion of electronic states within the valence and conduction bands. This challenging measurement has only been possible here thanks to sample rastering during data acquisition. Our intensity maps show experimental effective masses in good agreement with first-principles calculations. First, excited electrons relax to the bottom of the conduction band, and then they form bound states with the underlying holes. The wavevector tomography of the resulting wavefunction is compatible with Wannier excitons with a Bohr radius of 2.8 nm. No evidence of self-trapping is observed until 120 ps after photoexcitation. Moreover, a thorough analysis of the momentum distribution curve also excludes that fluctuating ferroelectric domains can give rise to dynamic Rashba splitting of very large amplitude.
@@
We acknowledge financial support of the 2D-HYPE project from the Agence Nationale de la Recherche (ANR, Nr. ANR 21-CE30-0059), of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation, Nr. 490867834), of the MINTOAURE project (ANR A-22-PETA-0015) and of the Région Ile de France via (FemtoARPES2.0 project N. EX079194). Weiyan Qi thanks the China Scholarship Council (CSC). Computational time was granted by GENCI (Project No. 544).